OpenStudy (anonymous):

Explain how to convert 96 oz to pounds.

OpenStudy (anonymous):

There are 16 oz in one pound. If we use dimensional analysis\[96 [oz] \cdot {1 \over 16} \left [ lb \over oz \right] = {96 \over 16} [lb]\]

OpenStudy (anonymous):

show work

OpenStudy (s):

1/16 = x/96 one pound / 16 ounces = x pounds / 96 ounces 96 = 16x cross multiply x = 96/16 divide x = 6 there are 6 pounds in 96 ounces